Fermi National Laboratory
 

Accelerator Update for December 29, 2008 - January 5, 2009

 

The Integrated Luminosity for 12/29/08 to 1/5/09 was 56.04 Pb-1.

Monday Dec 29

The day shift began with the Tevatron (TeV) down due to a sector A3 harmonic filter and VCB problem, with Operations monitoring the Recycler stash (387.4E10), with the Antiproton Source stacking (176.9E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store, and with the MiniBooNE and NuMI experiments taking beam.

During the day shift experts repaired the A3 harmonic filter. Operators turned on the TeV and checked the orbits. Pbar experts investigated a problem with an Accumulator RF station (ARF4). The TeV was put into shot setup by the end of the shift.

During the evening shift Operations established store 6692 with an initial luminosity of 346.53E30. Experts accessed the Booster and replaced a sump pump. Experts accessed MI and worked on a kicker magnet.

Tuesday Dec 30

The midnight shift began with Operations monitoring store 6692 (130.2E30) and the stash (186.2E10), with Pbar stacking (130.2E10), and with MiniBooNE and NuMI taking beam.

During the midnight shift a problem developed with the Pelletron VME. Controls experts investigated.

During the day shift operators terminated store 6692. A TeV flying wire got stuck in the beam at sector E11, which experts fixed. Operators put the TeV into shot setup.

During the evening shift Operations established store 6694 with an initial luminosity of 327.7E30.

Wednesday Dec 31

The midnight shift began with Operations monitoring store 6694 (127.67E30) and the stash (204.4E10), with Pbar stacking (23.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift a Pbar Accumulator RF station (ARF2) tripped off. It reset.

During the day shift Operations terminated store 6694 and established store 6695 with an initial luminosity of 347.05E30. MI experts conducted studies.

During the evening shift MIRF14 tripped off. It reset. The Pelletron VME failed again.

Thursday Jan 1

The midnight shift began with Operations monitoring store 6695 (84.32E30) and the stash (341.0E10), with Pbar stacking (30.3E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift Operations terminated store 6695 and established store 6696 with an initial luminosity of 352.3E30. The Pelletron tripped off due to a controls problem, which experts fixed.

During the day shift a kicker (KPS2A) fired early during a stash and kicked out ~110E10 of antiprotons. Linac suffered from a quadrupole power supply (QPS121) regulation problem.

During the evening shift kicker experts replaced a power supply and a cooling fan for KPS2A.

Friday Jan 2

The midnight shift began with Operations monitoring store 6696 and the stash (336.4E10), with Pbar stacking (33.1E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift Operations terminated store 6696 and put the TeV into shot setup, but problems with protons halted the setup. Booster and MI experts came in to work on machine efficiencies. MI experts investigated and found a problem with the dampers. A Recycler kicker (KPS3A) tripped off due to a low insulating gas (flourinert) level.

During the day shift MI experts found a blown fuse on a beam detector that was causing the damper problems. Operations established store 6698 with an initial luminosity of 305.92E30. Kicker experts accessed MI to fill KPS3A.

During the evening shift Kicker experts completed their work on KPS3A.

Saturday Jan 3

The midnight shift began with Operations monitoring the TeV store 6698 (125.65E30) and the stash (272.1E10), with Pbar stacking (18.6E10), and with MiniBooNE and NuMI taking beam.

During the midnight shift store 6698 aborted when the TeV sector C0 shunt power supply tripped off. TeV experts investigated, but could not find a cause for the trip. Operations put the TeV into shot setup.

During the day shift Operations established store 6701 with an initial luminosity of 313.3E30. A pulsed magnet LCW leak in the Pbar target station halted stacking. Experts accessed and tightened a fitting.

During the evening shift the Pbar pulsed magnet leak resumed and halted stacking. Operations terminated store 6701. Pbar experts accessed and replaced a fitting a on the Pbar water skid.

Sunday Jan 4

The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI taking beam.

During the midnight shift D0 experimenters accessed their collision hall to repair a power supply. Operations put the TeV into shot setup and established store 6704 with an initial luminosity of 308.69E30. The Pelletron VME tripped off.

During the day shift a Recycler beam valve failed and the stash was aborted. Experts accessed the MI tunnel and began repairs.

During the evening shift Recycler experts replaced a bad air pressure regulator for the beam valve. Operators replaced a Linac quadrupole (QPS117) power supply regulation card.

Monday Jan 5

The midnight shift began with Operations monitoring the store and stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift a MI transfer line Lambertson power supply (LAM52) tripped off. Power supply experts came in and repaired the supply.

The Plans for Monday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/05/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory